what comes next in the series 5 6 14 45 184 .....
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

what comes next in the series 5 6 14 45 184 .....
-------
Ans: 460
-----------------------
Using a Cubic Regression program I get
y = -0.4267x^3-54.38x^2+110.19x-60.2
---
f(6) = 460
===============
Cheers,
Stan H.
===============

Answer by Edwin McCravy(20056)

General term solutions are not unique..  The other tutor
used a regression formula with a graphing calculator.  But
the most logical pattern is:

 Start with 5
  5×1+1 =   6
  6×2+2 =  14
 14×3+3 =  45
 45×4+4 = 184,  so
184×5+5 = 925

Edwin